OpenStudy (anonymous):

Find the limit as x approaches 0 [(1/square root of 1+x)-1]/x

zepdrix (zepdrix):

\[\lim_{x \rightarrow 0} \frac{ \frac{ 1 }{ \sqrt{1+x} }-1 }{ x }\] Hmm looks like this?

OpenStudy (anonymous):

Yes it does

zepdrix (zepdrix):

What have you tried so far? There are a couple approaches, you can get a common denominator in the top, then multiply top/bottom by the CONJUGATE of the top. Or alternatively, you could multiply the top/bottom by the conjugate of the top right now, but it might be a little messy doing it that way :D

zepdrix (zepdrix):

Still confused lills? Need to see the steps? :o

OpenStudy (anonymous):

I started with timesing by radical 1+x but I got confused after that

zepdrix (zepdrix):

 Were you able to get to here ok? :D